\begin{thebibliography}{99}




%\cite{Rubin:1980zd}
\bibitem{Rubin:1980zd}
  V.~C.~Rubin, N.~Thonnard and W.~K.~Ford, Jr.,
  ``Rotational properties of 21 SC galaxies with a large range of luminosities and radii, from NGC 4605 /R = 4kpc/ to UGC 2885 /R = 122 kpc/,''
  Astrophys.\ J.\  {\bf 238} (1980) 471.
 % doi:10.1086/158003
  %%CITATION = doi:10.1086/158003;%%
  %520 citations counted in INSPIRE as of 18 Jun 2016


%\cite{Clowe:2006eq}
\bibitem{Clowe:2006eq}
  D.~Clowe, M.~Bradac, A.~H.~Gonzalez, M.~Markevitch, S.~W.~Randall, C.~Jones and D.~Zaritsky,
  ``A direct empirical proof of the existence of dark matter,''
  Astrophys.\ J.\  {\bf 648} (2006) L109
 % doi:10.1086/508162
  [astro-ph/0608407].
  %%CITATION = doi:10.1086/508162;%%
  %1591 citations counted in INSPIRE as of 27 Dec 2019

\bibitem{ghost}
S.~Danieli, P.~van~Dokkum, R.~Abraham, C.~Conroy, A.~E.~Dolphin and  A.~J.~Romanowsky, 
``A Tip of the Red Giant Branch Distance to the Dark Matter Deficient Galaxy NGC 1052-DF4 from Deep Hubble Space Telescope Data,"
 	arXiv:1910.07529 [astro-ph.GA].

%\cite{Milgrom:1983ca}
\bibitem{Milgrom:1983ca}
  M.~Milgrom,
  ``A Modification of the Newtonian dynamics as a possible alternative to the hidden mass hypothesis,''
  Astrophys.\ J.\  {\bf 270} (1983) 365.
  %doi:10.1086/161130
  %%CITATION = doi:10.1086/161130;%%
  %1340 citations counted in INSPIRE as of 18 Jun 2016





 
  
  
  
  








%\cite{Siegel:1993xq}
\bibitem{Siegel:1993xq}
  W.~Siegel,
  ``Two vierbein formalism for string inspired axionic gravity,''
  Phys.\ Rev.\ D {\bf 47} (1993) 5453 
  %doi:10.1103/PhysRevD.47.5453
  [hep-th/9302036].
  %%CITATION = doi:10.1103/PhysRevD.47.5453;%%
  %187 citations counted in INSPIRE as of 18 Jun 2016



%\cite{Siegel:1993th}
\bibitem{Siegel:1993th}
  W.~Siegel,
  ``Superspace duality in low-energy superstrings,''
  Phys.\ Rev.\ D {\bf 48} (1993) 2826 
 % doi:10.1103/PhysRevD.48.2826
  [hep-th/9305073].
  %%CITATION = doi:10.1103/PhysRevD.48.2826;%%
  %230 citations counted in INSPIRE as of 18 Jun 2016



%\cite{Hull:2009mi}
\bibitem{Hull:2009mi}
  C.~Hull and B.~Zwiebach,
  ``Double Field Theory,''
  JHEP {\bf 0909} (2009) 099 
  %doi:10.1088/1126-6708/2009/09/099
  [arXiv:0904.4664 [hep-th]].
  %%CITATION = doi:10.1088/1126-6708/2009/09/099;%%
  %267 citations counted in INSPIRE as of 18 Jun 2016



%\cite{Hull:2009zb}
\bibitem{Hull:2009zb}
  C.~Hull and B.~Zwiebach,
  ``The Gauge algebra of double field theory and Courant brackets,''
  JHEP {\bf 0909} (2009) 090 
 % doi:10.1088/1126-6708/2009/09/090
  [arXiv:0908.1792 [hep-th]].
  %%CITATION = doi:10.1088/1126-6708/2009/09/090;%%
  %184 citations counted in INSPIRE as of 18 Jun 2016



%\cite{Hohm:2010jy}
\bibitem{Hohm:2010jy}
  O.~Hohm, C.~Hull and B.~Zwiebach,
  ``Background independent action for double field theory,''
  JHEP {\bf 1007} (2010) 016 
  %doi:10.1007/JHEP07(2010)016
  [arXiv:1003.5027 [hep-th]].
  %%CITATION = doi:10.1007/JHEP07(2010)016;%%
  %216 citations counted in INSPIRE as of 18 Jun 2016



%\cite{Hohm:2010pp}
\bibitem{Hohm:2010pp}
  O.~Hohm, C.~Hull and B.~Zwiebach,
  ``Generalized metric formulation of double field theory,''
  JHEP {\bf 1008} (2010) 008 
  %doi:10.1007/JHEP08(2010)008
  [arXiv:1006.4823 [hep-th]].
  %%CITATION = doi:10.1007/JHEP08(2010)008;%%
  %223 citations counted in INSPIRE as of 18 Jun 2016



%\cite{Jeon:2010rw}
\bibitem{Jeon:2010rw}
  I.~Jeon, K.~Lee and J.~H.~Park,
  ``Differential geometry with a projection: Application to double field theory,''
  JHEP {\bf 1104} (2011) 014 
  %doi:10.1007/JHEP04(2011)014
  [arXiv:1011.1324 [hep-th]].
  %%CITATION = doi:10.1007/JHEP04(2011)014;%%
  %101 citations counted in INSPIRE as of 18 Jun 2016


%\cite{Jeon:2011cn}
\bibitem{Jeon:2011cn}
  I.~Jeon, K.~Lee and J.~H.~Park,
  ``Stringy differential geometry, beyond Riemann,''
  Phys.\ Rev.\ D {\bf 84} (2011) 044022
  % doi:10.1103/PhysRevD.84.044022
  [arXiv:1105.6294 [hep-th]].
  %%CITATION = doi:10.1103/PhysRevD.84.044022;%%
  %115 citations counted in INSPIRE as of 23 Mar 2017




%\cite{Park:2013mpa}
\bibitem{Park:2013mpa}
  J.~H.~Park,
  ``Comments on double field theory and diffeomorphisms,''
  JHEP {\bf 1306} (2013) 098
 % doi:10.1007/JHEP06(2013)098
  [arXiv:1304.5946 [hep-th]].
  %%CITATION = doi:10.1007/JHEP06(2013)098;%%
  %55 citations counted in INSPIRE as of 23 Mar 2017

%\cite{Park:2015bza}
\bibitem{Park:2015bza}
  J.~H.~Park, S.~J.~Rey, W.~Rim and Y.~Sakatani,
  ``$\ODD$ covariant Noether currents and global charges in double field theory,''
  JHEP {\bf 1511} (2015) 131
  % doi:10.1007/JHEP11(2015)131
  [arXiv:1507.07545 [hep-th]].
  %%CITATION = doi:10.1007/JHEP11(2015)131;%%
  %11 citations counted in INSPIRE as of 23 Mar 2017




%\cite{Angus:2018mep}
\bibitem{Angus:2018mep}
  S.~Angus, K.~Cho and J.~H.~Park,
  ``Einstein Double Field Equations,''
  Eur.\ Phys.\ J.\ C {\bf 78} (2018) no.6,  500
 % doi:10.1140/epjc/s10052-018-5982-y
  [arXiv:1804.00964 [hep-th]].
  %%CITATION = doi:10.1140/epjc/s10052-018-5982-y;%%
  %1 citations counted in INSPIRE as of 23 





%\cite{Jeon:2011kp}
\bibitem{Jeon:2011kp}
  I.~Jeon, K.~Lee and J.~H.~Park,
  ``Double field formulation of Yang--Mills theory,''
  Phys.\ Lett.\ B {\bf 701} (2011) 260
 % doi:10.1016/j.physletb.2011.05.051
  [arXiv:1102.0419 [hep-th]].
  %%CITATION = doi:10.1016/j.physletb.2011.05.051;%%
  %49 citations counted in INSPIRE as of 15 Jul 2019



%\cite{Hohm:2011ex}
\bibitem{Hohm:2011ex}
  O.~Hohm and S.~K.~Kwak,
  ``Double Field Theory Formulation of Heterotic Strings,''
  JHEP {\bf 1106} (2011) 096
  %doi:10.1007/JHEP06(2011)096
  [arXiv:1103.2136 [hep-th]].
  %%CITATION = doi:10.1007/JHEP06(2011)096;%%
  %114 citations counted in INSPIRE as of 15 Jul 2019


%\cite{Jeon:2011vx}
\bibitem{Jeon:2011vx}
  I.~Jeon, K.~Lee and J.~H.~Park,
  ``Incorporation of fermions into double field theory,''
  JHEP {\bf 1111} (2011) 025 
%  doi:10.1007/JHEP11(2011)025 
 [arXiv:1109.2035 [hep-th]].
  %%CITATION = doi:10.1007/JHEP11(2011)025;%%
  %53 citations counted in INSPIRE as of 18 Jun 2016


%\cite{Coimbra:2011nw}
\bibitem{Coimbra:2011nw}
  A.~Coimbra, C.~Strickland-Constable and D.~Waldram,
  ``Supergravity as Generalised Geometry I: Type II Theories,''
  JHEP {\bf 1111} (2011) 091
  %doi:10.1007/JHEP11(2011)091
  [arXiv:1107.1733 [hep-th]].
  %%CITATION = doi:10.1007/JHEP11(2011)091;%%
  %150 citations counted in INSPIRE as of 09 Jul 2017



  %\cite{Rocen:2010bk}
\bibitem{Rocen:2010bk}
  A.~Rocen and P.~West,
  ``E11, generalised space-time and IIA string theory: the R-R sector,''
  doi:10.1142/9789814412551$\underline{\,\,\,}$0020
  arXiv:1012.2744 [hep-th].
  %%CITATION = doi:10.1142/9789814412551_0020;%%
  %38 citations counted in INSPIRE as of 14 Oct 2019
  
  





  
%\cite{Hohm:2011zr}
\bibitem{Hohm:2011zr}
  O.~Hohm, S.~K.~Kwak and B.~Zwiebach,
  ``Unification of Type II Strings and T-duality,''
  Phys.\ Rev.\ Lett.\  {\bf 107} (2011) 171603
 % doi:10.1103/PhysRevLett.107.171603
  [arXiv:1106.5452 [hep-th]].
  %%CITATION = doi:10.1103/PhysRevLett.107.171603;%%
  %101 citations counted in INSPIRE as of 15 Jul 2019
    
 %\cite{Hohm:2011dv}
\bibitem{Hohm:2011dv}
  O.~Hohm, S.~K.~Kwak and B.~Zwiebach,
  ``Double Field Theory of Type II Strings,''
  JHEP {\bf 1109} (2011) 013
%  doi:10.1007/JHEP09(2011)013
  [arXiv:1107.0008 [hep-th]].
  %%CITATION = doi:10.1007/JHEP09(2011)013;%%
  %106 citations counted in INSPIRE as of 22 Feb 2018



%\cite{Jeon:2012kd}
\bibitem{Jeon:2012kd}
  I.~Jeon, K.~Lee and J.~H.~Park,
  ``Ramond-Ramond Cohomology and O(D,D) T-duality,''
  JHEP {\bf 1209} (2012) 079
%  doi:10.1007/JHEP09(2012)079
  [arXiv:1206.3478 [hep-th]].
  %%CITATION = doi:10.1007/JHEP09(2012)079;%%
  %66 citations counted in INSPIRE as of 15 Jul 2019



%\cite{Jeon:2011sq}
\bibitem{Jeon:2011sq}
  I.~Jeon, K.~Lee and J.~H.~Park,
  ``Supersymmetric Double Field Theory: Stringy Reformulation of Supergravity,''
  Phys.\ Rev.\ D {\bf 85} (2012) 081501(R), 
   Erratum: [Phys.\ Rev.\ D {\bf 86} (2012) 089903]
 % doi:10.1103/PhysRevD.86.089903, 10.1103/PhysRevD.85.081501, 10.1103/PhysRevD.85.089908
  [arXiv:1112.0069 [hep-th]].
  %%CITATION = doi:10.1103/PhysRevD.86.089903, 10.1103/PhysRevD.85.081501, 10.1103/PhysRevD.85.089908;%%
  %100 citations counted in INSPIRE as of 27 Dec 2019
  
  

  

%\cite{Jeon:2012hp}
\bibitem{Jeon:2012hp}
  I.~Jeon, K.~Lee, J.~H.~Park and Y.~Suh,
  ``Stringy Unification of Type IIA and IIB Supergravities under N=2 D=10 Supersymmetric Double Field Theory,''
  Phys.\ Lett.\ B {\bf 723} (2013) 245. 
  %doi:10.1016/j.physletb.2013.05.016
  [arXiv:1210.5078 [hep-th]].
  %%CITATION = doi:10.1016/j.physletb.2013.05.016;%%
  %44 citations counted in INSPIRE as of 18 Jun 2016




%\cite{Ko:2016dxa}
\bibitem{Ko:2016dxa}
  S.~M.~Ko, J.~H.~Park and M.~Suh,
  ``The rotation curve of a point particle in stringy gravity,''
  JCAP {\bf 1706} (2017) 002
%  doi:10.1088/1475-7516/2017/06/002
  [arXiv:1606.09307 [hep-th]].
  %%CITATION = doi:10.1088/1475-7516/2017/06/002;%%
  %15 citations counted in INSPIRE as of 22 Dec 2019

%\cite{Blair:2017gwn}
\bibitem{Blair:2017gwn}
  C.~D.~A.~Blair,
  ``Particle actions and brane tensions from double and exceptional geometry,''
  JHEP {\bf 1710} (2017) 004
 % doi:10.1007/JHEP10(2017)004
  [arXiv:1707.07572 [hep-th]].
  %%CITATION = doi:10.1007/JHEP10(2017)004;%%
  %10 citations counted in INSPIRE as of 27 Dec 2019





%\cite{Hull:2006va}
\bibitem{Hull:2006va}
  C.~M.~Hull,
  ``Doubled Geometry and T-Folds,''
  JHEP {\bf 0707} (2007) 080
 % doi:10.1088/1126-6708/2007/07/080
  [hep-th/0605149].
  %%CITATION = doi:10.1088/1126-6708/2007/07/080;%%
  %220 citations counted in INSPIRE as of 26 Aug 2018


%\cite{Hull:2009sg}
\bibitem{Hull:2009sg}
  C.~M.~Hull and R.~A.~Reid-Edwards,
  ``Non-geometric backgrounds, doubled geometry and generalised T-duality,''
  JHEP {\bf 0909} (2009) 014
  doi:10.1088/1126-6708/2009/09/014
  [arXiv:0902.4032 [hep-th]].
  %%CITATION = doi:10.1088/1126-6708/2009/09/014;%%
  %100 citations counted in INSPIRE as of 23 Oct 2019




%\cite{Lee:2013hma}
\bibitem{Lee:2013hma}
  K.~Lee and J.~H.~Park,
  ``Covariant action for a string in "doubled yet gauged" spacetime,''
  Nucl.\ Phys.\ B {\bf 880} (2014) 134  
  %doi:10.1016/j.nuclphysb.2014.01.003
   [arXiv:1307.8377 [hep-th]].
  %%CITATION = doi:10.1016/j.nuclphysb.2014.01.003;%%
  %28 citations counted in INSPIRE as of 18 Jun 2016



%\cite{Blair:2013noa}
\bibitem{Blair:2013noa}
  C.~D.~A.~Blair, E.~Malek and A.~J.~Routh,
  ``An $O(D, D)$ invariant Hamiltonian action for the superstring,''
  Class.\ Quant.\ Grav.\  {\bf 31} (2014) no.20,  205011
  %doi:10.1088/0264-9381/31/20/205011
  [arXiv:1308.4829 [hep-th]].
  %%CITATION = doi:10.1088/0264-9381/31/20/205011;%%
  %27 citations counted in INSPIRE as of 27 Dec 2019


%\cite{Park:2016sbw}
\bibitem{Park:2016sbw}
  J.~H.~Park,
  ``Green-Schwarz superstring on doubled-yet-gauged spacetime,''
  JHEP {\bf 1611} (2016) 005
  % doi:10.1007/JHEP11(2016)005
  [arXiv:1609.04265 [hep-th]].
  %%CITATION = doi:10.1007/JHEP11(2016)005;%%
  %2 citations counted in INSPIRE as of 23 Mar 2017

%\cite{Blair:2019qwi}
\bibitem{Blair:2019qwi}
  C.~D.~A.~Blair,
 ``A worldsheet supersymmetric Newton-Cartan string,''
  JHEP {\bf 1910} (2019) 266
%  doi:10.1007/JHEP10(2019)266
  [arXiv:1908.00074 [hep-th]].
  %%CITATION = doi:10.1007/JHEP10(2019)266;%%
  %4 citations counted in INSPIRE as of 27 Dec 2019


  %\cite{Choi:2015bga}
\bibitem{Choi:2015bga}
  K.~S.~Choi and J.~H.~Park,
  ``Standard Model as a Double Field Theory,''
  Phys.\ Rev.\ Lett.\  {\bf 115} (2015) no.17,  171603
 % doi:10.1103/PhysRevLett.115.171603
  [arXiv:1506.05277 [hep-th]].
  %%CITATION = doi:10.1103/PhysRevLett.115.171603;%%
  %10 citations counted in INSPIRE as of 16 Jul 2019




%\cite{Park:2019hbc}
\bibitem{Park:2019hbc}
  J.~H.~Park,
  ``$\mathbf{O}(D,D)$ completion of the Einstein Field Equations,''
  PoS CORFU {\bf 2018} (2019) 145
  doi:10.22323/1.347.0145
  [arXiv:1904.04705 [hep-th]].
  %%CITATION = doi:10.22323/1.347.0145;%%
  %3 citations counted in INSPIRE as of 27 Dec 2019




%\cite{Morand:2017fnv}
\bibitem{Morand:2017fnv}
  K.~Morand and J.~H.~Park,
  ``Classification of non-Riemannian doubled-yet-gauged spacetime,''
  Eur.\ Phys.\ J.\ C {\bf 77} (2017) no.10,  685
 % doi:10.1140/epjc/s10052-017-5257-z
  [arXiv:1707.03713 [hep-th]].
  %%CITATION = doi:10.1140/epjc/s10052-017-5257-z;%%
  %4 citations counted in INSPIRE as of 31 Dec 2017

%\cite{Cho:2018alk}
\bibitem{Cho:2018alk}
  K.~Cho, K.~Morand and J.~H.~Park,
  ``Kaluza--Klein reduction on a maximally non-Riemannian space is moduli-free,''
  Phys.\ Lett.\ B {\bf 793} (2019) 65
 % doi:10.1016/j.physletb.2019.04.042
  [arXiv:1808.10605 [hep-th]].
  %%CITATION = doi:10.1016/j.physletb.2019.04.042;%%
  %9 citations counted in INSPIRE as of 31 Dec 2019



%\cite{Berman:2019izh}
\bibitem{Berman:2019izh}
  D.~S.~Berman, C.~D.~A.~Blair and R.~Otsuki,
  ``Non-Riemannian geometry of M-theory,''
  JHEP {\bf 1907} (2019) 175
%  doi:10.1007/JHEP07(2019)175
  [arXiv:1902.01867 [hep-th]].
  %%CITATION = doi:10.1007/JHEP07(2019)175;%%
  %13 citations counted in INSPIRE as of 27 Dec 2019




%\cite{Cho:2019ofr}
\bibitem{Cho:2019ofr}
  K.~Cho and J.~H.~Park,
  ``Remarks on the non-Riemannian sector in Double Field Theory,''
  arXiv:1909.10711 [hep-th].
  %%CITATION = ARXIV:1909.10711;%%
  %2 citations counted in INSPIRE as of 23 Dec 2019



%\cite{Park:2017snt}
\bibitem{Park:2017snt}
  J.~H.~Park,
  ``Stringy Gravity: Solving the Dark Problems at ‘short’ distance,''
  EPJ Web Conf.\  {\bf 168} (2018) 01010
  %doi:10.1051/epjconf/201816801010
  [arXiv:1707.08961 [hep-th]].
  %%CITATION = doi:10.1051/epjconf/201816801010;%%


\bibitem{Cheung:2008fz}
  Y.~K.~E.~Cheung and F.~Xu,
  ``Fitting the Galaxy Rotation Curves: Strings versus NFW profile,''
  arXiv:0810.2382 [hep-th].
  %%CITATION = ARXIV:0810.2382;%%
  %6 citations counted in INSPIRE as of 22 Dec 2019




%\cite{Kim:1979if}
\bibitem{Kim:1979if}
  J.~E.~Kim,
  ``Weak Interaction Singlet and Strong CP Invariance,''
  Phys.\ Rev.\ Lett.\  {\bf 43} (1979) 103.
  %%CITATION = PRLTA,43,103;%%
  %1296 citations counted in INSPIRE as of 17 juin 2015
  
%\cite{Shifman:1979if}
\bibitem{Shifman:1979if}
  M.~A.~Shifman, A.~I.~Vainshtein and V.~I.~Zakharov,
  ``Can Confinement Ensure Natural CP Invariance of Strong Interactions?,''
  Nucl.\ Phys.\ B {\bf 166} (1980) 493.
  %%CITATION = NUPHA,B166,493;%%
  %1042 citations counted in INSPIRE as of 17 juin 2015
  
  
%\cite{Dine:1981rt}
\bibitem{Dine:1981rt}
  M.~Dine, W.~Fischler and M.~Srednicki,
  ``A Simple Solution to the Strong CP Problem with a Harmless Axion,''
  Phys.\ Lett.\ B {\bf 104} (1981) 199.
  %%CITATION = PHLTA,B104,199;%%
  %1583 citations counted in INSPIRE as of 17 juin 2015
  
%\cite{Zhitnitsky:1980tq}
\bibitem{Zhitnitsky:1980tq}
  A.~R.~Zhitnitsky,
  ``On Possible Suppression of the Axion Hadron Interactions. (In Russian),''
  Sov.\ J.\ Nucl.\ Phys.\  {\bf 31} (1980) 260
   [Yad.\ Fiz.\  {\bf 31} (1980) 497].
  %%CITATION = SJNCA,31,260;%%
  %798 citations counted in INSPIRE as of 17 juin 2015


%\cite{Svrcek:2006yi}
\bibitem{Svrcek:2006yi}
  P.~Svrcek and E.~Witten,
  ``Axions In String Theory,''
  JHEP {\bf 0606} (2006) 051
%  doi:10.1088/1126-6708/2006/06/051
  [hep-th/0605206].
  %%CITATION = doi:10.1088/1126-6708/2006/06/051;%%
  %754 citations counted in INSPIRE as of 30 Jan 2020     

  
  %\cite{Angus:2019bqs}
\bibitem{Angus:2019bqs}
  S.~Angus, K.~Cho, G.~Franzmann, S.~Mukohyama and J.~H.~Park,
  ``$\mathbf{O}(D,D)$ completion of the Friedmann equations,''
  arXiv:1905.03620 [hep-th].
  %%CITATION = ARXIV:1905.03620;%%
  %3 citations counted in INSPIRE as of 29 Dec 2019






%\cite{Jackson:1998nia}
\bibitem{Jackson:1998nia}
  J.~D.~Jackson,
  ``Classical Electrodynamics,'' 3rd ed. 1998,    832 Pages,  
Wiley, ISBN: 978-0-471-30932-1.
  %%CITATION = INSPIRE-1390077;%%
  %247 citations counted in INSPIRE as of 23 Dec 2019

  
  
 
%\cite{Burgess:1994kq}
\bibitem{Burgess:1994kq}
  C.~P.~Burgess, R.~C.~Myers and F.~Quevedo,
  ``On spherically symmetric string solutions in four-dimensions,''
  Nucl.\ Phys.\ B {\bf 442} (1995) 75
 % doi:10.1016/S0550-3213(95)00090-9
  [hep-th/9410142].
  %%CITATION = doi:10.1016/S0550-3213(95)00090-9;%%
  %21 citations counted in INSPIRE as of 09 Feb 2018



%\cite{Hernquist:1990be}
\bibitem{Hernquist:1990be}
  L.~Hernquist,
  ``An Analytical Model for Spherical Galaxies and Bulges,''
  Astrophys.\ J.\  {\bf 356} (1990) 359.
%  doi:10.1086/168845
  %%CITATION = doi:10.1086/168845;%%
  %1535 citations counted in INSPIRE as of 22 Dec 2019

%\cite{Navarro:1995iw}
\bibitem{Navarro:1995iw}
  J.~F.~Navarro, C.~S.~Frenk and S.~D.~M.~White,
  ``The Structure of cold dark matter halos,''
  Astrophys.\ J.\  {\bf 462} (1996) 563
%  doi:10.1086/177173
  [astro-ph/9508025].
  %%CITATION = doi:10.1086/177173;%%
  %4654 citations counted in INSPIRE as of 22 Dec 2019


 %\cite{Andringa:2012uz}
\bibitem{Andringa:2012uz}
  R.~Andringa, E.~Bergshoeff, J.~Gomis and M.~de Roo,
  ``'Stringy' Newton-Cartan Gravity,''
  Class.\ Quant.\ Grav.\  {\bf 29} (2012) 235020
 % doi:10.1088/0264-9381/29/23/235020
  [arXiv:1206.5176 [hep-th]].
  %%CITATION = doi:10.1088/0264-9381/29/23/235020;%%
  %55 citations counted in INSPIRE as of 30 Dec 2019

%\cite{Hansen:2018ofj}
\bibitem{Hansen:2018ofj}
  D.~Hansen, J.~Hartong and N.~A.~Obers,
  ``Action Principle for Newtonian Gravity,''
  Phys.\ Rev.\ Lett.\  {\bf 122} (2019) no.6,  061106
 % doi:10.1103/PhysRevLett.122.061106
  [arXiv:1807.04765 [hep-th]].
  %%CITATION = doi:10.1103/PhysRevLett.122.061106;%%
  %33 citations counted in INSPIRE as of 30 Dec 2019

%\cite{Bergshoeff:2019pij}
\bibitem{Bergshoeff:2019pij}
  E.~A.~Bergshoeff, J.~Gomis, J.~Rosseel, C.~Simsek and Z.~Yan,
  ``String Theory and String Newton-Cartan Geometry,''
  J.\ Phys.\ A {\bf 53} (2020) no.1,  014001
 % doi:10.1088/1751-8121/ab56e9
  [arXiv:1907.10668 [hep-th]].
  %%CITATION = doi:10.1088/1751-8121/ab56e9;%%
  %8 citations counted in INSPIRE as of 30 Dec 2019


%\cite{Hansen:2019vqf}
\bibitem{Hansen:2019vqf}
  D.~Hansen, J.~Hartong and N.~A.~Obers,
  ``Gravity between Newton and Einstein,''
  Int.\ J.\ Mod.\ Phys.\ D {\bf 28} (2019) no.14,  1944010
  %doi:10.1142/S0218271819440103
  [arXiv:1904.05706 [gr-qc]].
  %%CITATION = doi:10.1142/S0218271819440103;%%
  %12 citations counted in INSPIRE as of 30 Dec 2019

%\cite{Basile:2019pic}
\bibitem{Basile:2019pic}
  T.~Basile, E.~Joung and J.~H.~Park,
  ``A note on Faddeev--Popov action for doubled-yet-gauged particle and graded Poisson geometry,''
  arXiv:1910.13120 [hep-th].
  %%CITATION = ARXIV:1910.13120;%%
  
  
 \bibitem{Eikonal}
 E. J. Post, 
`` Formal Structure of Electromagnetics,"  1997, 224 pages, 
Dover Publications,  ISBN-13: 978-0486654270.
  
\end{thebibliography}

\documentclass[twocolumn,showpacs,preprintnumbers,amsmath,amssymb]{revtex4-1}
\usepackage{graphicx}% Include figure files
\usepackage{dcolumn}% Align table columns on decimal point
\usepackage{bm}% bold math
\usepackage[compact]{titlesec}
\usepackage{hyperref}
\newcommand\bmtimes{{\bm{\times}}}
\newcommand\bmcdot{{\bm{\cdot}}}

\newcommand\ret{{\mathrm{ret}}}

\newcommand\To{T_{\scriptscriptstyle{{(0)}}}}
\newcommand\brTo{\brT_{\scriptscriptstyle{{(0)}}}}

\newcommand{\half}{{{\textstyle\frac{1}{2}}}}
\newcommand{\quarter}{{{\textstyle\frac{1}{4}}}}
\newcommand{\be}{\begin{equation}}
\newcommand{\ee}{\end{equation} }
\newcommand{\beqa}{\begin{eqnarray} }
\newcommand{\eeqa}{\end{eqnarray} }
\newcommand{\ba}{\begin{array}}
\newcommand{\ea}{\end{array}}
\newcommand{\bpm}{\begin{pmatrix}}
\newcommand{\epm}{\end{pmatrix}}

\newcommand{\ODD}{\mathbf{O}(D,D)}

\newcommand\Phis{\Phi_{{\bfs}}}
\newcommand\EK{K}

\newcommand\rd{{\rm d}}

\newcommand\bfx{{\mathbf{x}}{}}
\newcommand\bfs{{\mathbf{s}}{}}

\newcommand\bfB{{\mathbf{B}}}

\newcommand\bfH{{\mathbf{H}}}
\newcommand\bfK{{\mathbf{K}}}

\newcommand\bfS{{\mathbf{S}}}

\newcommand\eff{{\rm{\bf{eff.}}}}
\newcommand\rhoeff{{\rho_{\eff}}\!}

\newcommand\cH{{\cal H}}

\newcommand\bfhn{{\mathbf{\hat{n}}}}

\newcommand\whK{{\widehat{K}}}
\newcommand\whT{{\widehat{T}}}

\newcommand\dis{\displaystyle}

\def\brh{\bar{h}}

\def\brB{\bar{B}}

\def\brH{\bar{H}}
\def\brK{\bar{K}}

\def\brT{{\bar{T}}}

\newcommand{\brphi}{\bar{\phi}}

\newcommand{\na}{{\nabla}}
\newcommand{\bmna}{\bm{\nabla}}
\newcommand{\trd}{{\bigtriangledown}}

\begin{document}

%\preprint{APS/123-QED}
\title{Stringy Newton Gravity with $H$-flux}
\author{Kyoungho Cho}
\email{khcho23@sogang.ac.kr}
\author{Kevin Morand}
\email{morand@sogang.ac.kr}
\author{Jeong-Hyuck Park\,}
\email{park@sogang.ac.kr}

\affiliation{\vspace{1pt}Department of Physics, Sogang University, 35 Baekbeom-ro, Mapo-gu, Seoul 04107,  Korea}

\begin{abstract}
\centering\begin{minipage}{\dimexpr\paperwidth-6.7cm}
\noindent  A Symmetry Principle   has  been shown to   augment  unambiguously  the  Einstein Field Equations,  promoting         the whole    closed-string massless NS-NS sector  to stringy graviton fields. Here we consider   its  weak field approximation,  take a non-relativistic limit, and   derive the stringy   augmentation of  Newton Gravity:  
\vspace{-12pt}
\[
\begin{array}{lll}
\bm{\nabla}^{2}\Phi=4\pi G \rho+\mathbf{H}{\bm{\cdot}}\mathbf{H}\,,
\quad&\qquad\bm{\nabla}{\bm{\cdot}}\mathbf{H}=0\,,
\quad&\qquad
\bm{\nabla}\bm{\times}\mathbf{H}=4\pi G\, \mathbf{K}\,.
\end{array}\vspace{-7pt}
\]
Not only the mass density~$\rho$ but also the   current density $\mathbf{K}$ is  intrinsic to matter.  Sourcing  $\mathbf{H}$ which is of NS-NS $H$-flux origin,   $\mathbf{K}$ is    nontrivial  if the matter is   `stringy'. $\mathbf{H}$ contributes quadratically  to the Newton potential, but otherwise is decoupled from  the  point particle dynamics, \textit{i.e.~}$\mathbf{\ddot{x}}=-\bm{\nabla}\Phi$. 
We define `stringization'  analogous to magnetization  and  discuss regular as well as  monopole-like singular solutions. 
\end{minipage}
\end{abstract}

%\pacs{11.25.-w}% PACS, the Physics and Astronomy
                             % Classification Scheme.
                             
%%%
%%11.25.-w 	Strings and branes                             
%\keywords{Suggested keywords}%Use showkeys class option if keyword
                              %display desired
\maketitle
\section{Introduction}
\vspace{5pt}
One of the fundamental  problems  in physics today  is the dark matter problem. Despite  a variety of   observational indications,  \textit{e.g.~}galaxy rotation curves~\cite{Rubin:1980zd},  the Bullet Cluster (1E0657-558)~\cite{Clowe:2006eq}, and ghostly galaxies without  dark matter~\cite{ghost},    no single experiment  has ever  succeeded  in the direct detection of  dark matter particles.    Alternative hypotheses,  including notably  MOND~\cite{Milgrom:1983ca},  pass over the    Equivalence Principle  and   modify General Relativity (GR),     but only give  partial  explanations while  being often   accused  of harming the mathematical beauty thereof.% from an aesthetic perspective.

Developments   over the last decade in an area of string theory, now called  Double Field Theory (DFT),  have gradually  unveiled  a new  form of pure gravity~\cite{Siegel:1993xq,Siegel:1993th,Hull:2009mi,Hull:2009zb,Hohm:2010jy,
Hohm:2010pp,Jeon:2010rw,Jeon:2011cn,Park:2013mpa,Park:2015bza,
Angus:2018mep}. It is based on the  $\ODD$ Symmetry Principle, with   $D$ denoting   the spacetime dimension.  The symmetry can be broken only spontaneously but never explicitly, as the theory is constructed in terms of strictly  $\ODD$-covariant  field variables, namely   the DFT-dilaton~$d$ and DFT-metric~$\cH_{AB}$ (or more powerfully DFT-vielbeins),  forming  the new pure gravity sector.  The Symmetry Principle   further fixes    their  coupling  to generic matter contents which should   also be  in $\ODD$ representations. Examples include  Yang--Mills~\cite{Jeon:2011kp,Hohm:2011ex},  fermions~\cite{Jeon:2011vx} (\textit{c.f.~}\cite{Coimbra:2011nw}), R-R sector~\cite{Rocen:2010bk,Hohm:2011zr,Hohm:2011dv,Jeon:2012kd}, full-order supersymmetrizations~\cite{Jeon:2011sq,Jeon:2012hp}, point particles~\cite{Ko:2016dxa,Blair:2017gwn},
fundamental strings~\cite{Hull:2006va,Hull:2009sg,Lee:2013hma,
Blair:2013noa,Park:2016sbw,Blair:2019qwi}, and the Standard Model itself~\cite{Choi:2015bga}. Naturally,  the Einstein Field Equations are   augmented to an $\ODD$-symmetric form~\cite{Angus:2018mep} (\textit{c.f.~}\cite{Park:2019hbc} for a short summary):
\be
G_{AB}={\frac{8\pi G}{c^{4}}} T_{AB}\,,
\label{MASTER}
\ee
which carry $\ODD$ vector indices and unify the Euler--Lagrange equations of all  the stringy graviton fields, 
$\left\{\cH_{AB},d\right\}$.  In a parallel manner to GR,  $G_{AB}$ is the off-shell conserved $\ODD$-symmetric Einstein curvature   constructed out of $\left\{\cH_{AB},d\right\}$~\cite{Park:2015bza}, while $T_{AB}$ is the on-shell conserved energy-momentum  tensor, defined through the variation of the matter Lagrangian with respect to 
$\left\{\cH_{AB},d\right\}$~\cite{Angus:2018mep}.

Remarkably, the perfectly $\ODD$-symmetric vacua, satisfying ${G_{AB}=0}$, turned out to  be a  topological phase  which allows    no moduli and  no  interpretation within   Riemannian geometry, thus  escaping   beyond the realm of GR~\cite{Lee:2013hma,Morand:2017fnv,Cho:2018alk}. Only after a spontaneous symmetry breaking of $\ODD$, the familiar string theory backgrounds characterized  by the Riemannian metric $g_{\mu\nu}$ and the Kalb--Ramond skew-symmetric two-form  potential $B_{\mu\nu}$ emerge:  these component fields  parametrize  the DFT-metric while  being identified   as the Nambu--Goldstone bosons~\cite{Berman:2019izh}. The master formula~(\ref{MASTER}) then 
reduces to  (\textit{c.f.~}\cite{Cho:2019ofr} for non-Riemannian cases)
\vspace{-5pt}
\be
\ba{r}
\half e^{2\phi}\trd^{\rho}\!\left(e^{-2\phi}H_{\rho\mu\nu}\right)=\,\frac{8\pi G}{c^{4}}\EK_{[\mu\nu]}\,,\\
R_{\mu\nu}+2\trd_{\mu}(\partial_{\nu}\phi)-\quarter H_{\mu\rho\sigma}H_{\nu}{}^{\rho\sigma}
=\frac{8\pi G}{c^{4}}\EK_{(\mu\nu)}\,,\\
R+4\trd_{\mu}(\partial^{\mu}\phi)-4\partial_{\mu}\phi\partial^{\mu}\phi-\textstyle{\frac{1}{12}}H_{\lambda\mu\nu}H^{\lambda\mu\nu}=\,\frac{8\pi G}{c^{4}}\To\,,~~
\ea
\label{REDFE}
\ee
which imply a pair of conservation laws, \vspace{-5pt} 
\be
\ba{r}
\na^{\mu}\!\left(e^{-2\phi}\EK_{[\mu\nu]}\right)=0\,,\\
\na^{\mu}\EK_{(\mu\nu)}-2\partial^{\mu}\phi\,\EK_{(\mu\nu)}+\half H_{\nu}{}^{\lambda\mu}\EK_{[\lambda\mu]}-\half\partial_{\nu}\To=0\,.
\ea
\label{CON}
\ee
The left and the right  hand sides of the equalities in (\ref{REDFE})  come from $G_{AB}$ and  $T_{AB}$ in (\ref{MASTER}) separately. Schematically, $K_{[\mu\nu]}$, $K_{(\mu\nu)}$, and $\To$  are   the energy-momentum tensor   components relevant to $B_{\mu\nu}$,  $g_{\mu\nu}$,  and the string dilaton $\phi=d+\quarter\ln\left|g\right|$, respectively.  Having said that,  since (\ref{MASTER}) is derived from the variations of the $\ODD$ covariant fields, in particular the DFT-dilaton $d$ rather than the string dilaton $\phi$, the  specific form of the  Einstein curvature, $R_{\mu\nu}-\half g_{\mu\nu}R$,  does not automatically appear  in (\ref{REDFE}).

We stress that for each  matter content, the $\ODD$ symmetry  determines   its  coupling to the closed-string massless NS-NS sector of $\{B_{\mu\nu},g_{\mu\nu},\phi\}$ and hence fixes  $K_{\mu\nu},\To$ completely~\cite{Angus:2018mep}.  For example,  a point particle should couple to the string frame metric~$g_{\mu\nu}$ only   ---minimally in the standard way--- resulting in  $K_{[\mu\nu]}=0$, $\To=0$, and 
\be
\dis{
K^{\mu\nu}(x)=\frac{mc}{2}\int\rd\tau~\dot{y}^{\mu}(\tau)\dot{y}^{\nu}(\tau)\frac{\delta\big(x-y(\tau)\big)e^{2\phi}}{\sqrt{-g}}\,.}
\label{Kparticle}
\ee
The particle  follows    geodesics defined in the string frame,
\be
\ddot{x}^{\lambda}+\gamma^{\lambda}_{\mu\nu}\dot{x}^{\mu}\dot{x}^{\nu}=0\,.
\label{geodesic}
\ee
Rewriting this  in the Einstein frame would involve  the gradient of $\phi$ and thus obscure the Equivalence Principle. That is to say, the $\ODD$ Symmetry Principle asserts  that  the Equivalence Principle  for a point particle should  hold    in the string frame. On the other hand, fundamental strings, spinorial fermions, and  the R-R sector couple  to $B_{\mu\nu}$ and $g_{\mu\nu}$ but not to  $\phi$, resulting in  asymmetric $K_{\mu\nu}$ and (still) trivial $\To$.  Further,  in contrast,  gauge bosons   couple to  $g$ and $\phi$, but do not interact with the $B$-field~(\ref{photon})\,\cite{Choi:2015bga}. While the $B$-field  does not interact with the electromagnetic force nor point particles,  its electric $H$-flux  nevertheless contributes to the mass formula~\cite{Angus:2018mep},
\be
M=\int e^{-2d}\left|\,2K_{0}{}^{0}+\frac{1}{16\pi G}H_{0\mu\nu}H^{0\mu\nu}\,\right|\,.
\label{MASS}
\ee
In fact, having a larger profile than the (visible) matter represented  here by $K_{0}{}^{0}$, the electric $H$-flux can produce non-monotonic non-Keplerian rotation curves~\cite{Ko:2016dxa}.  In this way,   $H$-flux  behaves  like   dark matter~\cite{Park:2017snt}, \textit{c.f.~}\cite{Cheung:2008fz}.  Note that its dual scalar  is also known as  a dark matter candidate,  `an axion'~\cite{Kim:1979if,Shifman:1979if,Dine:1981rt,Zhitnitsky:1980tq,Svrcek:2006yi}.

It is the purpose of the present paper to consider the weak field approximation of (\ref{REDFE}), (\ref{CON}), (\ref{geodesic}) for  ${D=4}$,   take a consistent non-relativistic limit,  and  faithfully derive the stringy augmentation of Newton Gravity   spelled out  in the Abstract. We hope our work may deepen  the  physical   understanding of    the $\ODD$-completed General Relativity~(\ref{MASTER}) and contribute to   examining rigorously   the prospect of  $H$-flux as a dark matter candidate.  

%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%
%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%

\vspace{5pt}
\section{Lorentz symmetric weak field approximation\label{SEClEDFE}}
\vspace{5pt}
We  start our  weak field approximation of (\ref{REDFE}) by linearizing  the metric, 
$g_{\mu\nu}=\eta_{\mu\nu}+h_{\mu\nu}$, $g^{\mu\nu}\simeq\eta^{\mu\nu}-h^{\mu\nu}$, around a flat Minkowskian background  with trivial $H$-flux and  dilaton~$\phi$. The spacetime indices are raised or lowered by the constant   metric~$\eta$, \textit{e.g.~}$h^{\mu\nu}=\eta^{\mu\rho}\eta^{\nu\sigma}h_{\rho\sigma}$.
The linearized  Riemann curvature,
\be
R^{\kappa}{}_{\lambda\mu\nu}=\half\left(
\partial_{\lambda}\partial_{\mu}h^{\kappa}{}_{\nu}
-\partial_{\lambda}\partial_{\nu}h^{\kappa}{}_{\mu}
-\partial^{\kappa}\partial_{\mu}h_{\lambda\nu}
+\partial^{\kappa}\partial_{\nu}h_{\lambda\mu}\right)\,,\vspace{-3pt}
\label{lR}
\ee
is invariant under the linearized diffeomorphisms,\vspace{-8pt}
\be
\ba{ll}
\delta_{\xi} h_{\mu\nu}=\partial_{\mu}\xi_{\nu}+\partial_{\nu}\xi_{\mu}\,,\qquad&\qquad
\delta_{\xi} R^{\kappa}{}_{\lambda\mu\nu}= 0\,.
\ea
\ee
Using $\delta_{\xi}(\partial_{\rho}h^{\rho}{}_{\mu}-\half\partial_{\mu}h^{\rho}{}_{\rho})=\partial_{\rho}\partial^{\rho}\xi_{\mu}$, we fix the gauge,
\be
\partial_{\rho}h^{\rho}{}_{\mu}-\half\partial_{\mu}h^{\rho}{}_{\rho}+2\partial_{\mu}\phi=0\,.
\label{gauge}
\ee
It follows from (\ref{lR}) and (\ref{gauge}),\vspace{-8pt}
\be
\ba{r}
R_{\mu\nu}= -\half\partial_{\rho}\partial^{\rho}h_{\mu\nu}-2\partial_{\mu}\partial_{\nu}\phi\,,~\\
\partial_{\rho}\partial^{\rho}\phi=\quarter \partial_{\rho}\partial^{\rho}h^{\sigma}{}_{\sigma}-\half\partial_{\rho}\partial_{\sigma}h^{\rho\sigma}\,,
\ea
\ee
and further, from the integrability of (\ref{gauge}),
\be
\partial_{\mu}\partial_{\rho}h^{\rho}{}_{\nu}=\partial_{\nu}\partial_{\rho}h^{\rho}{}_{\mu}\,.
\label{integrability}
\ee
Similarly, we choose a  gauge  for the $B$-field,\vspace{-8pt}
\be
\ba{lll}
\delta_{\lambda}B_{\mu\nu}=\partial_{\mu}\lambda_{\nu}-\partial_{\nu}\lambda_{\mu}\quad&\quad\Longrightarrow\quad&\quad
\partial_{\rho}B^{\rho}{}_{\mu}=0\,.
\ea\label{gaugeB}
\ee
Now, assuming the following  scales, 
\be
h_{\mu\nu}\,\sim\, K_{(\mu\nu)}\,\sim\, \phi\,\sim\, \To\,\sim\, \left(H_{\lambda\mu\nu}\right)^{2}\,\sim\,\left(K_{[\mu\nu]}\right)^{2}\,,
\label{lscale}
\ee
the consistent  linearization of (\ref{REDFE})  can be achieved,  \vspace{-8pt}
\be
\ba{rcl}
\partial^{\rho}H_{\rho\mu\nu}\,=\,\partial_{\rho}\partial^{\rho}B_{\mu\nu}&=&\frac{16\pi G}{c^{4}}\EK_{[\mu\nu]}\,,\\
\partial_{\rho}\partial^{\rho}h_{\mu\nu}+\half H_{\mu\rho\sigma}H_{\nu}{}^{\rho\sigma}&=&-\frac{16\pi G}{c^{4}}\EK_{(\mu\nu)}\,,\\
\partial_{\rho}\partial_{\sigma}h^{\rho\sigma}+\textstyle{\frac{1}{12}}H_{\rho\sigma\tau}H^{\rho\sigma\tau}&=&-\frac{8\pi G}{c^{4}}\To\,,
\ea
\label{lEDFE}
\ee
which, with  (\ref{integrability}),  imply   the following linearized conservation equations, \vspace{-8pt}
\be
\ba{ll}
\partial^{\rho}K_{[\rho\mu]}=0\,,\quad&\quad
\partial^{\rho}K_{(\rho\mu)}+\half H_{\mu}{}^{\rho\sigma}K_{[\rho\sigma]}-\half\partial_{\mu}\To=0\,.
\ea
\label{lCON}
\ee
With a well-known  relation,
\be
\!\!\bmna^{2}\frac{1}{\left|\bfx-\bfx^{\prime}\right|}=
-\bmna\bmcdot\left(\frac{\bfx-\bfx^{\prime}}{\,\,\left|\bfx-\bfx^{\prime}\right|^{3}}\right)
=-4\pi\delta(\bfx-\bfx^{\prime})\,,
\label{standard}
\ee
the first formula in (\ref{lEDFE})  is solved by  a  `retarded potential', 
\be
B_{\mu\nu}(x)=-\dis{\frac{4G}{c^{4}}\int\rd^{3}x^{\prime}~\frac{K_{[\mu\nu]}(x^{0}_{\ret}\,,\,\bfx^{\prime})}{\left|\bfx-\bfx^{\prime}\right|}}\,,
\label{Bmunu}
\ee
where $x^{0}_{\ret}:=x^{0}-\left|\bfx-\bfx^{\prime}\right|$ is the   retarded temporal coordinate.  Thanks to $\partial^{\rho}K_{[\rho\mu]}=0$ (\ref{lCON}), up to a surface integral,  the gauge condition~(\ref{gaugeB}) is indeed fulfilled, 
\be
\dis{
\partial^{\mu}B_{\mu\nu}(x)=-\frac{4G}{c^{4}}\int\rd^{3}x^{\prime}~\frac{\left.\,\partial^{\prime\mu}K_{[\mu\nu]}(x^{\prime})\right|_{x^{\prime 0}\rightarrow x^{0}_{\ret}}}{\left|\bfx-\bfx^{\prime}\right|}=0\,.}
\label{gaugeB2}
\ee
The expression for the $H$-flux follows 
\be
H_{\lambda\mu\nu}(x)=-\dis{\frac{4G}{c^{4}}\int\rd^{3}x^{\prime}~\frac{\,\left.3\partial^{\prime}_{[\lambda}K_{\mu\nu]}(x^{\prime})\right|_{x^{\prime 0}\rightarrow x^{0}_{\ret}}}{\left|\bfx-\bfx^{\prime}\right|}}\,.
\label{Hlmn}
\ee
In a similar fashion,  with some shorthand notations,  
\be
\ba{l}
\whK_{(\mu\nu)}:=\frac{16\pi G}{c^{4}}\EK_{(\mu\nu)}+\half H_{\mu\rho\sigma}H_{\nu}{}^{\rho\sigma}\,,\\
\whT:=\frac{8\pi G}{c^{4}}\To+\textstyle{\frac{1}{12}}H_{\rho\sigma\tau}H^{\rho\sigma\tau}\,,
\ea
\label{barred}
\ee  
we solve the second formula in (\ref{lEDFE}), 
\be
h_{\mu\nu}(x)=\dis{\frac{1}{4\pi}\int\rd^{3}x^{\prime}~\frac{\whK_{(\mu\nu)}(x^{0}_{\ret}\,,\,\bfx^{\prime})}{\left|\bfx-\bfx^{\prime}\right|}}\,.
\label{hmunu}
\ee 
From (\ref{lEDFE}), (\ref{lCON}), and the $H$-flux Bianchi identity,  we get
\be
\dis{\partial^{\mu}\whK_{(\mu\nu)}=\partial_{\nu}\whT}\,,
\ee
which gives,  comparable  to (\ref{gaugeB2}),  
\be
\partial_{\rho}h^{\rho}{}_{\mu}(x)
=\dis{\partial_{\mu}\left[\,\frac{1}{4\pi}\int\rd^{3}x^{\prime}~\frac{\whT(x^{0}_{\ret}\,,\,\bfx^{\prime})}{\left|\bfx-\bfx^{\prime}\right|}\,\right]}\,.
\label{pmu}
\ee
This verifies  for consistency  that (\ref{hmunu})  indeed satisfies   the integrability condition~(\ref{integrability}), while    
 the third formula in (\ref{lEDFE}) is automatically fulfilled as $\partial_{\rho}\partial_{\sigma}h^{\rho\sigma}(x)=-\whT(x)$. 
Lastly, from (\ref{gauge}),  the string dilaton is fixed, 
\be
\phi(x)=\frac{1}{4\pi}\dis{\int\rd^{3}x^{\prime}~
\left.\frac{\,\frac{1}{4}\whK_{\mu}{}^{\mu}(x^{\prime})-\half\whT(x^{\prime})\,}
{\left|\bfx-\bfx^{\prime}\right|}\right|_{x^{\prime 0}\,\rightarrow\, x^{0}_{\ret}}}\,.
\label{phifixed}
\ee
To summarize, through (\ref{Hlmn}), (\ref{hmunu}), and (\ref{phifixed}),   $\{K_{\mu\nu},\To\}$ (`matter') determines $\{H_{\lambda\mu\nu},h_{\mu\nu},\phi\}$ (`geometry').   While $H_{\lambda\mu\nu}$ is given by a single volume integral,  the other two, $h_{\mu\nu},\phi$, involve triple volume integrals  of `matter'.  It is  worth while to note that the  quantity inside the bracket in (\ref{pmu}) is related to the $\ODD$ singlet integral measure of DFT (scalar density with weight one),
\be
e^{-2d}\!=\!\sqrt{-g}e^{-2\phi}\simeq 1+\half h_{\mu}{}^{\mu}-2\phi=1+\textstyle{\frac{1}{4\pi}
}\dis{\int}\textstyle{\!\rd^{3}x^{\prime}\,\frac{\whT(x^{0}_{\ret}\,,\,\bfx^{\prime})}{\left|\bfx-\bfx^{\prime}\right|}\,.}
\label{em2d}
\ee
The linearized geodesic equation assumes  the form,
\be
\ddot{x}^{\lambda}+\half\!\left(\partial_{\mu}h^{\lambda}{}_{\nu}+\partial_{\nu}h^{\lambda}{}_{\mu}-\partial^{\lambda}h_{\mu\nu}\right)\dot{x}^{\mu}\dot{x}^{\nu}=0\,.
\label{lgeodesic}
\ee

%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%
%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%
%\vspace{5pt}
\section{Non-relativistic limit: Stringy  Newton \label{SECNONR}}
\vspace{5pt}
\noindent We proceed to take the non-relativistic,  large $c$ limit of (\ref{lEDFE}). With $x^{0}=ct$, we let all the fields be functions of $(t,\bfx)=(t, x^{i})$, $i=1,2,3$. We   suppress   $\partial_{0}{\,=\frac{1}{c}\frac{\partial~}{\partial t}}{\,= 0}$ and put  $\partial_{\rho}\partial^{\rho}{\,= \partial_{i}\partial^{i}}{\,=\bmna^{2}}$. Further, for the point particle of (\ref{Kparticle}), (\ref{geodesic}), we set  ${t=\tau}$ (proper time)   and $\left|\dot{\bfx}\right|<<c$, such that the linearized geodesic equation~(\ref{lgeodesic}) reduces  to $\ddot{x}^{i}=\half \partial_{i}h_{00}c^{2}$ only.  Combined with (\ref{gauge}),  (\ref{lEDFE}),  this  fixes all the scales, which are  compatible  with  (\ref{lscale}): \vspace{-7pt}
\be
\ba{rlrl}
h_{\mu\nu}&\!=c^{-2}\brh_{\mu\nu}\,,\quad&\qquad  K_{(\mu\nu)}&\!=c^{2}\brK_{(\mu\nu)}\,,\\
\phi&\!=c^{-2}\brphi\,,\quad&\qquad \To&\!=c^{2}\brTo\,,\\
B_{\mu\nu}&\!=c^{-1}\brB_{\mu\nu}\,,\quad&\qquad K_{[\mu\nu]}&\!=c^{3}\brK_{[\mu\nu]}\,,\\
H_{\lambda\mu\nu}&\!=c^{-1}\brH_{\lambda\mu\nu}\,.\quad&\quad{}&{}
\ea
\label{nrscales}
\ee
This then    straightforwardly  produces  the non-relativistic limit of the previous  weak field approximation:  with  ${\partial_{0}=0}$ assumed, after removing the symbol~$c$, replacing   $x^{0}_{\ret}$ by $t$,   and putting  bars over  each  quantity,  all the formulas  from section~\ref{SEClEDFE}  survive to preserve    their forms and set of  relations.      For example, (\ref{lEDFE}) reduces  to\vspace{-5pt}
\be
\ba{rcl}
\partial^{i}\brH_{i\mu\nu}\,=\,\bmna^{2}\brB_{\mu\nu}&=&16\pi G\brK_{[\mu\nu]}\,,\\
\bmna^{2}\brh_{\mu\nu}+\half \brH_{\mu\rho\sigma}\brH_{\nu}{}^{\rho\sigma}
&=&-16\pi G\brK_{(\mu\nu)}\,,\\
\partial_{i}\partial_{j}\brh^{ij}+\textstyle{\frac{1}{12}}\brH_{\rho\sigma\tau}\brH^{\rho\sigma\tau}&=&-8\pi G\brTo\,,
\ea
\label{nREDFE}
\ee
%and the gauge  conditions~(\ref{gauge}), (\ref{integrability}), (\ref{gaugeB}) give
%\be
%\ba{lll}
%\partial_{i}\brh^{i}{}_{0}=0\,,\quad&\quad
%\partial_{i}\partial_{k}\brh^{k}{}_{j}=
%\partial_{j}\partial_{k}\brh^{k}{}_{i}\,,\quad&\quad\partial^{i}\brB_{i\mu}=0\,.
%\ea
%\label{nrgauge}
%\ee
%Setting the  scale of   $K_{[\mu\nu]}$ one power higher than  $K_{(\mu\nu)}$ in (\ref{nrscales}), we  are able to keep the stringy property of matter, if any,  and let the $H$-flux survive in the limit. 
%From the consevation relations~(\ref{lCON}), the matter content should meet
%\be
%\ba{cc}
%\multicolumn{2}{c}{
%\partial^{j}\brK_{(ji)}+\half %\brH_{ijk}\brK^{[jk]}+\brH_{ij0}\brK^{[j0]}-\half\partial_{i}\brTo=0\,,}\\
%\partial^{j}\brK_{(j0)}+\half \brH_{0jk}\brK^{[jk]}=0\,,\qquad&\qquad
%\partial^{i}\brK_{[i\mu]}=0\,.
%\ea
%\label{nRCON}
%\ee
and (\ref{Bmunu})  becomes\vspace{-5pt}
\be
\brB_{\mu\nu}(x)=-\dis{4G\int\rd^{3}x^{\prime}~\frac{\brK_{[\mu\nu]}(t\,,\,\bfx^{\prime})}{\left|\bfx-\bfx^{\prime}\right|}}\,.
\label{nrBmunu}
\ee
Furthermore,  the  $H$-flux Bianchi identity reads  now 
\be
\partial_{[i}\brH_{jk]0}=0\,.
\label{nrBianchi}
\ee
Hereafter, we focus on the  Newton potential which is the only quantity directly relevant to the particle dynamics, \vspace{-7pt}
\be
\ba{ll}
\Phi:=-\half c^{2}h_{00}=-\half\brh_{00}\,,\quad&\quad\ddot{\bfx}=-\bmna\Phi\,.
\ea
\ee
We then identify all the quantities which can affect the  Newton potential: namely,  
 the mass density~$\rho$, the stringy current density~$\bfK$, and $B$-field/$H$-flux vectors~$\bfB,\bfH$, as follows\vspace{-5pt}
\be
\ba{l}
\rho:=2\brK_{00}\,,\\
\bfK:=2\sqrt{2}\left(\brK_{[01]},\,\brK_{[02]},\,\brK_{[03]}\right)\,,\\
\bfB:=\frac{1}{\sqrt{2}}\left(\brB_{10},\,\brB_{20},\,\brB_{30}\right)\,,\\
\bfH:=\bmna\bmtimes\bfB=\frac{1}{\sqrt{2}}\left(\brH_{023},\,\brH_{031},\,\brH_{012}\right)\,.
\ea
\ee
Crucially, $\{\rho,\bfK,\Phi,\bfH\}$ forms an `autonomy' of closed  relations: from (\ref{nREDFE}), \vspace{-5pt}
\be
\ba{ll}
\bmna^{2}\Phi=4\pi G \rho+\bfH\bmcdot\bfH\,,\qquad&\qquad
\bmna\bmtimes\bfH=4\pi G\, \bfK\,,
\ea
\label{MAIN}
\ee
and, from (\ref{lCON}), (\ref{nrBianchi}),  $\bfK$ and $\bfH$ are both divergenceless, \vspace{-5pt}
\be
\ba{ll}
\bmna\bmcdot\bfK=0\,,\qquad&\qquad\bmna\bmcdot\bfH=0\,.\vspace{-3pt}
\ea
\label{divergenceless}
\ee 
That is to say, the Newton potential  is fully determined by both  the mass density and the stringy  current density: directly so by $\rho$, and   indirectly  so by $\bfK$ as  mediated through     $\bfH$,
\be
\ba{l}
\dis{
\Phi=-G\!\int\!\rd^{3}x^{\prime}\,\frac{\, \rhoeff(t,\bfx^{\prime})\,}{\left|\bfx-\bfx^{\prime}\right|}\,,}\quad~~~~~\,\,
\rhoeff:=\rho+\textstyle{\frac{1}{4\pi G}}\bfH\bmcdot\bfH\,,\\
\dis{
\bfH=
G\!\int\!\rd^{3}x^{\prime}~\bfK(t,\bfx^{\prime})\bmtimes
\frac{(\bfx-\bfx^{\prime})\,}{\left|\bfx-\bfx^{\prime}\right|^{3}}=\bmna\bmtimes\bfB}\,,\\
\dis{
\bfB=G\!\int\!\rd^{3}x^{\prime}~~
\frac{\bfK(t,\bfx^{\prime})}{\left|\bfx-\bfx^{\prime}\right|}\,,\qquad~\,\,~~~\,~\bmna\bmcdot\bfB=0\,.}
\ea
\label{NewtonPot}
\ee
Although $\bfH$ originates from the electric components of the $H$-flux,  its behaviour is identical to the  magnetic field in classical Magnetostatics  such that  a  ``Biot--Savart law"  holds above. The vector potential $\bfB$  has  NS-NS $B$-field origin, and notation-wise should not  be confused with the magnetic field.    It is instructive to note that $\rho_{\eff}$ is consistent with (\ref{MASS}) at the linearized level (\textit{c.f.}~\cite{Angus:2019bqs}), and involves a double volume integral as
\be
\bfH\bmcdot\bfH=G^{2}\!
\int\!\rd^{3}x^{\prime}\!\int\!\rd^{3}x^{\prime\prime}~
\frac{\mathbf{{det}}{\tiny{\left[\ba{cc}(\bfx-\bfx^{\prime})\bmcdot
(\bfx-\bfx^{\prime\prime})&(\bfx-\bfx^{\prime})\bmcdot\bfK^{\prime\prime}\\
(\bfx-\bfx^{\prime\prime})\bmcdot\bfK^{\prime}&
\bfK^{\prime}\bmcdot\bfK^{\prime\prime}\ea\right]}}}
{\left|\bfx-\bfx^{\prime}\right|^{3}\left|\bfx-\bfx^{\prime\prime}\right|^{3}}\,,
\label{exact}
\ee
where  $\bfK^{\prime}=\bfK(t,\bfx^{\prime}), \,\bfK^{\prime\prime}=\bfK(t,\bfx^{\prime\prime})$.  \vspace{5pt}

In  analogy  to  the  magnetization in electrodynamics, 
 we introduce the notion of \textit{stringization}  for the  stringy  current density $\bfK$ which is divergence free, 
\be
\bfK(t,\bfx)=\bmna\bmtimes\bfs(t,\bfx)\,.
\ee
The corresponding  $\bfB$, $\bfH$ are, from (\ref{standard}), (\ref{NewtonPot})~(\textit{c.f.~}\cite{Jackson:1998nia}),
\be
\ba{lll}
\bfB&=&%\dis{G\!\int\!\rd^{3}x^{\prime}~\frac{\,\bmna^{\prime}\bmtimes\bfs(t,\bfx^{\prime})}{\left|\bfx-\bfx^{\prime}\right|}}\vspace{2pt}\\
%{}&=&
\dis{G\!\int\!\rd^{3}x^{\prime}~\frac{\bfs(t,\bfx^{\prime})\bmtimes(\bfx-\bfx^{\prime})}{\left|\bfx-\bfx^{\prime}\right|^{3}}+G\!\oint\rd\mathbf{A}\bmtimes\frac{\bfs(t,\bfx^{\prime})}{\left|\bfx-\bfx^{\prime}\right|}}\,,\vspace{10pt}\\
\bfH&=&\dis{4\pi G\bfs(t,\bfx)+G\!\int\!\rd^{3}x^{\prime}~\frac{3\bfhn^{\prime}\left(\bfhn^{\prime}\bmcdot\,\bfs(t,\bfx^{\prime})\right)-\bfs(t,\bfx^{\prime})}{\left|\bfx-\bfx^{\prime}\right|^{3}}}\vspace{-3pt}\\
{}&=&4\pi G\bfs(t,\bfx)- G\bmna\Phis(t,\bfx)\,,
\ea
\label{HPhis}
\ee
in which $\bfhn^{\prime}=\frac{\bfx-\bfx^{\prime}}{\left|\bfx-\bfx^{\prime}\right|}$   and $\Phis$ is a \textit{stringy scalar potential},
\be
\dis{
\Phis(t,\bfx)=\!\int\!\rd^{3}x^{\prime}\,\frac{\,\bfs(t,\bfx^{\prime})\bmcdot(\bfx-\bfx^{\prime})}{\left|\bfx-\bfx^{\prime}\right|^{3}}=-\bmna\bmcdot\!\!
\int\!\rd^{3}x^{\prime}\,\frac{\bfs(t,\bfx^{\prime})}{\left|\bfx-\bfx^{\prime}\right|}
\,.
}
\ee
Clearly from (\ref{HPhis}),   $\bmna\bmtimes\bfH=4\pi G\bmna\bmtimes\bfs$.  Far away from a localized  source, $\left|\bfx\right|{>>}\left|\bfx^{\prime}\right|$,  we observe a stringy dipole,   
\be
\ba{ll}
\dis{\bfH\simeq G\,\frac{3\hat{\bfx}\left(\hat{\bfx}\bmcdot\bfS(t)\right)-\bfS(t)}{\left|\bfx\right|^{3}}\,,}\quad&\quad
\bfS(t)%=\half\!\dis{\int}\!\rd^{3}x~\bfx\bmtimes\bfK(t,\bfx)
=\dis{\int}\!\rd^{3}x~\bfs(t,\bfx)\,.
\ea
\ee

As an example, we consider a uniformly  `stringized' sphere of radius $a$,  with   constant $\rho$ and $\bfs$, to get (\textit{c.f.~}\cite{Jackson:1998nia})
\be
\ba{lll}
\Phis=\frac{4\pi}{3}\,\bfs\bmcdot\bfx\,,\quad&\quad\!\!
\bfH=\frac{8\pi G}{3}\,\bfs\quad&~~\mbox{for~}\left|\bfx\right|\leq a\,,\\
\Phis=\frac{4\pi a^{3}}{3}\,\frac{\bfs\bmcdot\bfx}{\,\left|\bfx\right|^{3}}\,,\quad&\quad\!\!
\bfH=\frac{4\pi G a^{3}}{3}\left(\frac{3\hat{\bfx}(\hat{\bfx}\bmcdot\bfs)-\bfs}{\left|\bfx\right|^{3}}\right)\quad&~~\mbox{for~}\left|\bfx\right|> a\,.
\ea
\ee
Thus, the total effective  mass density~(\ref{NewtonPot}) reads
\be
\rhoeff(t,\bfx)=\left\{\ba{ll}
\rho+\frac{\,16\pi G }{9}\left|\bfs\right|^{2}
\quad&\quad\mbox{for~}\left|\bfx\right|\leq a\\
\frac{\,4\pi G}{9}\left|\bfs\right|^{2}a^{6}\!\left(\frac{1+3\cos^{2}\!\theta}{\left|\bfx\right|^{6}}\right)
\quad&\quad\mbox{for~}\left|\bfx\right|>a\,,
\ea
\right.
\label{rhoeff}
\ee
where $\theta$ is the angle between $\bfs$ and $\bfx$. As anticipated from the general formula~(\ref{exact}), $\rhoeff$ has a $\left|\bfx\right|^{-6}$ profile or `halo'.

Another example is  of Dirac monopole type (singular),
\be
\ba{ll}
\dis{\bfB=Gq\int_{\bfx^{\prime}=\bf{0}}^{\infty}\,\rd\bfx^{\prime}\bmtimes
\frac{(\bfx-\bfx^{\prime})}{\,\left|\bfx-\bfx^{\prime}\right|^{3}}\,,}\quad&~~
\dis{\bfH=Gq\frac{\bfx}{~\left|\bfx\right|^{3}}\,,}
\ea
\ee
where the path should not cross the point of $\bfx$. The profile is now thicker   as  $\rhoeff=\frac{Gq^{2}}{4\pi}\left|\bfx\right|^{-4}$ which may be  comparable to some known    dark matter profiles~\cite{Hernquist:1990be,Navarro:1995iw}.  In fact,  this configuration corresponds to the  linearization of   a known  exact spherical solution to (\ref{REDFE})~\cite{Burgess:1994kq} which has been shown to feature a non-monotonic and hence non-Keplerian rotation curve~\cite{Ko:2016dxa}.

%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%
%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%
\vspace{5pt}
\section{Conclusion\label{SECDISCUSSION}}
\vspace{5pt}
We have shown that  the $\ODD$-completed General Relativity or Einstein Field Equations~(\ref{MASTER}) reduce in the non-relativistic limit to the Stringy  Newton Gravity~(\ref{MAIN}). 
 Symmetry-wise, the $\ODD$ of (\ref{MASTER}) is broken spontaneously  in (\ref{REDFE})  of which  General Covariance is  reduced to Lorentz symmetry in (\ref{lEDFE}) and further to Galilean symmetry in (\ref{MAIN}). It would be of interest to investigate whether    General Covariance can be recovered  as in  (stringy) Newton--Cartan Gravity~\cite{Andringa:2012uz,Hansen:2018ofj,Bergshoeff:2019pij}.
 
 The  final resulting formulas~(\ref{MAIN})   resemble  a hybrid of  Newton Gravity and Magnetostatics. Out of $D^{2}+1=17$ number of  components of the stringy energy-momentum tensor, $T_{AB}$~(\ref{MASTER})   or $\{K_{\mu\nu},\To\}$~(\ref{REDFE}), only four, \textit{i.e.~}the   mass density~$\rho$ and  the stringy current density~$\bfK$, participate in determining the Newton potential~(\ref{NewtonPot}). Different types of matter have different $\rho$ and $\bfK$ (or \textit{stringization} $\bfs$).  If the matter is  non-stringy  particle-like, $\bfK$ is trivial and we fully recover  Newton Gravity. On the other hand,  in the presence of distinct  kinds of matter, the center of $\rho$ may not  coincide with  that of $\bfH\bmcdot\bfH$. These might explain the  ghostly galaxies without dark matter (or without  $H$-flux sourced by $\bfK$)~\cite{ghost}  and  the Bullet Cluster~\cite{Clowe:2006eq} respectively.   It would be of the utmost interest to test    Stringy  Newton Gravity more rigorously against observations. For this, one needs to also   analyze light  or electromagnetism (for a recent discussion see \cite{Hansen:2019vqf}).  The $\ODD$ Symmetry Principle prescribes   gauge bosons  to   couple to  $g$ and $\phi$  but  not to $B$-field~\cite{Choi:2015bga},
\be
S_{{\rm{photon}}}=\dis{\int\rd^{4}x~-\quarter\sqrt{-g}e^{-2\phi}g^{\mu\rho}g^{\nu\sigma}F_{\mu\nu}F_{\rho\sigma}\,.}
\label{photon}
\ee
This seems to imply that a photon would not merely follow a null geodesic, but would  be also  influenced by  the dilaton~$\phi$, \textit{e.g.~}\cite{Basile:2019pic}.  Further analysis on the action~(\ref{photon}), using methods  such as Eikonal approximation~\cite{Eikonal}, is desirable. \\
~\\
\indent\textit{Acknowledgments.}   We wish to thank Stephen Angus, David Berman,  Gabriele Gionti, Chethan Krishnan, and  Jae-Weon Lee for helpful discussions.  This work  was  supported by  the National Research Foundation of Korea   through  the Grants,  NRF-2016R1D1A1B01015196  and NRF-2018H1D3A1A01030137~(Korea Research Fellowship Program).
\hfill

\end{document}